OpenStudy (anonymous):

The vector position of a particle varies in time according to the expression r(vector)= 9.00 i(with arrow above)− 7.60t2 j(with arrow above) where r(vector) is in meters and t is in seconds. Calculate the particle's position and velocity at t = 2.00 s. r(vector)=___m

OpenStudy (anonymous):

the answer will be an equation

OpenStudy (anonymous):

r = -30.4tj

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

thats the vector i got that right

OpenStudy (anonymous):

i need the r(vector)

OpenStudy (anonymous):

v = dr/dt = 0 i + -15.2t j. v= -15.2(2)j =-30.4j \[r =\int\limits_{}^{}v.dt=\int\limits_{}^{}-30.4 .dt=-30.4tj\]

OpenStudy (anonymous):

i got that for v tho..its for r as well?

OpenStudy (anonymous):

its wrong